OpenStudy (vera_ewing):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the second equation is substituted into the first equation. 2x – 3y = –29 x + 4y = 13

OpenStudy (vera_ewing):

@amistre64 Can you check my answer? I think it's A. 2(4y + 13) – 3y = –29

OpenStudy (vera_ewing):

@tkhunny Is my answer correct?

OpenStudy (tkhunny):

No, sorry. x + 4y = 13 ==> x = 13 - 4y

OpenStudy (vera_ewing):

2(–4y + 13) – 3y = –29 would that be the answer? @tkhunny

OpenStudy (tkhunny):

You tell me. Did you substitute correctly? If so, then that is correct.

OpenStudy (vera_ewing):

I'm not sure if I did it right. can you just check to make sure you got the same thing as me?

OpenStudy (tkhunny):

There isn't anything else to do or to check. You're done. Why do you doubt?

OpenStudy (vera_ewing):

I just want to get it right. Did you get the same answer as me?

OpenStudy (tkhunny):

Asked and answered. Move on to the next one.